# Disentangled Representation Learning through Geometry Preservation with the Gromov-Monge Gap

### v1

@@ Abstract
Learning disentangled representations in an unsupervised manner is a fundamental challenge in machine learning. Solving it may unlock other problems, such as generalization, interpretability, or fairness. While remarkably difficult to solve in general, recent works have shown that disentanglement is provably achievable under additional assumptions that can leverage geometrical constraints, such as local isometry. To use these insights, we propose a novel perspective on disentangled representation learning built on quadratic optimal transport. Specifically, we formulate the problem in the Gromov-Monge setting, which seeks isometric mappings between distributions supported on different spaces. We propose the Gromov-Monge-Gap (GMG), a regularizer that quantifies the geometry-preservation of an arbitrary push-forward map between two distributions supported on different spaces. We demonstrate the effectiveness of GMG regularization for disentanglement on four standard benchmarks. Moreover, we show that geometry preservation can even encourage unsupervised disentanglement without the standard reconstruction objective - making the underlying model decoder-free, and promising a more practically viable and scalable perspective on unsupervised disentanglement.
@@ 1 Introduction
Learning low-dimensional representations of high-dimensional data is a fundamental challenge in unsupervised deep learning [5, 43], where the emphasis is put on learning representations that allow for efficient adaptation across a wide range of tasks [5, 28, 43]. Disentanglement [5, 27, 28, 43, 60] has shown significant promise in facilitating such generalization [5, 27, 28, 43, 44, 29, 60, 30, 4], alongside interpretability and fairness [42, 67]. Most works [5, 27, 33, 13, 43, 60] regard disentanglement as a one-to-one map between learned representations and ground-truth latent factors; seeking to recover these factors from data alone in an unsupervised fashion.
While unsupervised disentanglement is theoretically impossible [43], the inductive biases of autoencoder architectures ensure effective disentanglement in practice [59, 74, 29]. Most approaches operate on variational autoencoder (VAE) frameworks [35], using objectives that match latent VAE posteriors to factorized priors [27, 33, 38, 10, 13]. Recent works [29, 51, 31] provide a new perspective, showing how geometric constraints on representation spaces may enable disentanglement. In particular, Horan et al. [29] demonstrate that unsupervised disentanglement is always possible under the assumption of local isometry to the data and sufficient non-Gaussianity of the generative factors, thus supporting the isometry desideratum.
In this work, we show how these geometric desiderata can be effectively quantified through the lens of optimal transport (OT) theory [61, 55], by treating mapping to or from the latent space as transport maps $T$ from or to the data manifold, respectively. While the more classical tools described in the OT toolbox [55] provide tools to align distributions supported on the same space $\mathcal{X}$, using an inter-domain cost $c(\mathbf{x},\mathbf{y})$ for any two points $\mathbf{x},\mathbf{y}\in\mathcal{X}$, such approaches are not well defined when mapping between latent and data space with very different dimensionalities, for which no “natural” cost function exists.
Instead, we rely in this work on the Gromov-Wasserstein (GW) formalism [64, 48, 69, 63] formulation of OT, which considers instead intra-domain costs $c_{\mathcal{X}},c_{\mathcal{Y}}$ in each space. This formulation seeks instead the most geometry-preserving mapping between two distributions by favoring maps that minimize the distortion between intra-domain costs by comparing $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime}),c_{\mathcal{Y}}(T(\mathbf{x}),T(\mathbf{x}^{\prime}))$. While this distortion itself can be used as a regularization for a given map $T$ as done in Nakagawa et al. [51], we borrow a page from the Monge gap formalism proposed by Uscidda and Cuturi [68] to propose a modified and better-posed regularizer that substracts to the distortion of $T$, the best distortion that can be achieved by any map (or a proxy thereof).
In a nod to [68], we propose the novel Gromov-Monge gap (GMG), which measures if $T$ maps points while preserving geometric properties as much as possible - such as (scaled) isometry (distance preserving) or conformity (angle preserving). In contrast to the distortion, the GMG takes the most geometry-preserving mapping into account, resulting in a regularizer that is properly defined, as shown in our various results, showing notably that the GMG and its finite sample version are weakly convex. We show how the GMG can serve as an effective regularizer to different geometry-preserving desiderata [29, 41]. Motivated by this formalism, we build upon the Monge gap - a regularizer introduced in Uscidda and Cuturi [68] that measures whether a map $T$ transports a reference distribution at minimal displacement cost - to propose the novel Gromov-Monge gap (GMG), which measures if $T$ maps points while preserving geometric properties as much as possible - such as (scaled) isometry (distance preserving) or conformity (angle preserving). In contrast to the distortion, the GMG takes the most geometry-preserving mapping into account, resulting in a regularizer that is properly defined, as shown in our various results highlighting that the GMG and its finite sample version are weakly convex. We lay out how the GMG can serve as an effective regularizer to different geometry-preserving desiderata [29, 41].
Our experiments on four standard disentangled representation learning benchmarks show that the integration of these geometry-preserving desiderata through the Gromov-Monge Gap (GMG) significantly improves disentanglement performance across various methods, from the standard $\beta$-VAE to the combination of $\beta$-TCVAE with support factorization [60], outperforming a distortion-based regularization. Moreover, we demonstrate that these geometric regularizations can replace the standard reconstruction loss, enabling measurable unsupervised disentanglement even without a decoder, which is not feasible in standard frameworks that rely on the decoder-based reconstruction term to prevent collapse. This finding suggests the potential for more scalable unsupervised disentangled representation learning approaches and bridges to popular, weakly- or self-supervised encoder-only representation learning methods [14, 72, 3, 24].
@@ 2.1 Disentangled Representation Learning
Disentanglement has varying operational definitions [28, 43, 60]. In this work, we follow the common understanding [43, 44, 67, 60] where data $\mathbf{x}$ is generated by a process $p(\mathbf{x}|\mathbf{z})$ operating on ground-truth latent factors $\mathbf{z}\sim p(\mathbf{z})$, modeling underlying source of variations (s.a. object shape, color, background…). Given a dataset $\mathcal{D}=\{\mathbf{x}_{i}\}_{i=1}^{N}$, unsupervised disentangled representation learning aims to find a mapping $e_{\phi}$ s.t. $e_{\phi}(\mathbf{x}_{i})\approx\mathbb{E}[\mathbf{z}|\mathbf{x}_{i}]$, up to element-wise transformations. Notably, this is to be achieved without prior information on $p(\mathbf{z})$ and $p(\mathbf{x}|\mathbf{z})$.
@@
Most unsupervised disentanglement methods operate on variational autoencoders (VAEs)[35], which define a generative model of the form $p_{\theta}(\mathbf{x},\mathbf{z})=p(\mathbf{z})p_{\theta}(\mathbf{x}|\mathbf{z})$. Here, $p_{\theta}(\mathbf{x}|\mathbf{z})$ is a product of exponential family distributions with parameters computed by a decoder $d_{\theta}(\mathbf{z})$. The latent prior $p(\mathbf{z})$ is usually chosen to be a normal Gaussian $\mathcal{N}(\mathbf{0},\mathbf{I})$, and the probabilistic encoder $q_{\phi}(\mathbf{z}|\mathbf{x})$ is realized through a neural network $e_{\phi}(\mathbf{x})$ that predicts the parameters of the latent such that $q_{\phi}(\mathbf{z}|\mathbf{x})=\mathcal{N}(\mathbf{z}|e_{\phi}(\mathbf{x}))$. The $\beta$-VAE [27]
$\mathcal{L}_{\beta}(\theta,\phi)\vcentcolon=\mathbb{E}_{\mathbf{x}\sim p_{\textrm{data}},\mathbf{z}\sim q_{\phi}(\mathbf{z}|\mathbf{x})}[\log p_{\theta}(\mathbf{x}|\mathbf{z})-\beta D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z}))]$ (1)
achieves disentanglement by enforcing stronger, $\beta$-weighted prior matching on top of the reconstruction objective, assuming statistical factor independence [60]. Several follow-ups refine latent prior matching through different objectives or prior choices [13, 38, 10, 59].
@@
Recent studies [26, 12, 41, 51, 31] indicate that disentanglement can arise by encouraging learned representations to preserve meaningful geometric features of the data, such as scaled distances between samples. Notably, Horan et al. [29] demonstrated that disentanglement is provably feasible when the generative factors are sufficiently non-Gaussian and locally isometric to the data. In this work, we explore how to promote geometry preservation using quadratic OT between the latent and data spaces, which we introduce in the next section.
@@ 2.2 Quadratic Optimal Transport
OT [55] involves transferring one probability distribution to another while incorporating inductive biases. When these distributions lie on incomparable domains, the task is addressed using the Gromov-Monge and GW problems, also known as OT quadratic formulations. Formally, consider two compact $\mathcal{X}\subset\mathbb{R}^{d_{\mathcal{X}}}$, $\mathcal{Y}\subset\mathbb{R}^{d_{\mathcal{Y}}}$, each of them equipped with an intra-domain cost $c_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ and $c_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$. For $p\in\mathcal{P}(\mathcal{X})$ and $q\in\mathcal{P}(\mathcal{Y})$—two distributions supported on each domain—the Gromov-Monge problem [50] seeks a map $T:\mathcal{X}\to\mathcal{Y}$ that push-forwards $p$ onto $q$, while minimizing the distortion of the costs:
@@
When it exists, we call a solution $T^{\star}$ to (GMP) a Gromov-Monge map for costs $c_{\mathcal{X}},c_{\mathcal{Y}}$. However, solving this problem is difficult, and existence is not guaranteed in general [19]. Moreover, this formulation is ill-suited for discrete distributions $p,q$, as the constraint set might be empty. Replacing maps by coupling $\pi\in\Pi(p,q)$, i.e. distributions on $\mathcal{X}\times\mathcal{Y}$ with marginals $p$ and $q$, we obtain the Gromov-Wasserstein (GW) problem [48, 64]
$\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(p,q):=\min_{\pi\in\Pi(p,q)}\int_{(\mathcal{X}\times\mathcal{Y})^{2}}\tfrac{1}{2}|c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})|^{2}\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})\mathop{}\!\mathrm{d}\pi(\mathbf{x}^{\prime},\mathbf{y}^{\prime})\,.$ (GWP)
A solution $\pi^{\star}$ of (GWP) always exists, making $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}(p,q)$ a well-defined quantity. It quantifies the minimal distortion of the geometries induced by $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ achievable when coupling $p$ and $q$.
@@
When both $p$ and $q$ are instantiated as samples, the GW Prob. ((GWP)) translates to a quadratic assignment problem, whose objective can be regularized using entropy [15]. For empirical measures ${p}_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}}$, ${q}_{n}=\frac{1}{n}\sum_{j=1}^{n}\delta_{\mathbf{y}_{j}}$ and $\varepsilon\geq 0$, we set:
$\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{\varepsilon}({p}_{n},{q}_{n})\vcentcolon=\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{j})-c_{\mathcal{Y}}(\mathbf{y}_{i},\mathbf{y}_{j}))^{2}\mathbf{P}_{ij}\mathbf{P}_{i^{\prime}j^{\prime}}-\varepsilon H(\mathbf{P})\,,$ (EGWP)
where $U_{n}=\{\mathbf{P}\in\mathbb{R}^{n\times n}_{+},\mathbf{P}\mathbf{1}_{n}=\mathbf{P}^{T}\mathbf{1}_{n}=\tfrac{1}{n}\mathbf{1}_{n}\}$ and $H(\mathbf{P})=-\sum_{i,j=1}^{n}\mathbf{P}_{ij}\log(\mathbf{P}_{ij})$. As $\varepsilon\to 0$, we recover $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{0}=\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}$. In addition to yielding better statistical [73] and regularity [58] properties, entropic regularization also enhances computational performance. In practice, we can solve (EGWP) using a mirror descent scheme that iterates the Sinkhorn algorithm [54, 62].
Neural Solvers.
While for classical OT, numerous neural methods have been proposed [45, 37, 22, 68, 66], the GW setting has received less attention. To our knowledge, the only neural Gromov-Monge formulation proposed thus far is [52], which involves a min-max-min optimization procedure. On the other hand, Klein et al. [36] recently proposed an approach to compute neural GW couplings.

### v2

# Disentangled Representation Learning with the Gromov-Monge Gap

@@ Abstract
Learning disentangled representations from unlabelled data is a fundamental challenge in machine learning. Solving it may unlock other problems, such as generalization, interpretability, or fairness. Although remarkably challenging to solve in theory, disentanglement is often achieved in practice through prior matching. Furthermore, recent works have shown that prior matching approaches can be enhanced by leveraging geometrical considerations, e.g., by learning representations that preserve geometric features of the data, such as distances or angles between points. However, matching the prior while preserving geometric features is challenging, as a mapping that fully preserves these features while aligning the data distribution with the prior does not exist in general. To address these challenges, we introduce a novel approach to disentangled representation learning based on quadratic optimal transport. We formulate the problem using Gromov-Monge maps that transport one distribution onto another with minimal distortion of predefined geometric features, preserving them as much as can be achieved. To compute such maps, we propose the Gromov-Monge-Gap (GMG), a regularizer quantifying whether a map moves a reference distribution with minimal geometry distortion. We demonstrate the effectiveness of our approach for disentanglement across four standard benchmarks, outperforming other methods leveraging geometric considerations.
@@ 1 Introduction
Learning low-dimensional representations of high-dimensional data is a fundamental challenge in unsupervised deep learning (Bengio et al., 2014). Emphasis is put on learning representations that allow for efficient and robust adaptation across a wide range of tasks (Higgins et al., 2018, Locatello et al., 2019a). The fundamental property of disentanglement has shown significant promise to improve generalization (Locatello et al., 2020, Roth et al., 2023, Hsu et al., 2023, Barin-Pacela et al., 2024), interpretability and fairness (Locatello et al., 2019b, Träuble et al., 2021). Most works regard disentanglement as a one-to-one map between learned representations and ground-truth latent factors, effectively seeking to recover these factors from data alone in an unsupervised fashion. While unsupervised disentanglement is theoretically impossible (Locatello et al., 2019a), the inductive biases of autoencoder architectures ensure effective disentanglement in practice (Rolinek et al., 2019, Zietlow et al., 2021). Most approaches operate using variational autoencoder (VAE) frameworks (Kingma and Welling, 2014), using objectives that match latent VAE posteriors to factorized priors (Higgins et al., 2017, Kim and Mnih, 2018, Kumar et al., 2018, Burgess et al., 2018, Chen et al., 2018).
More recently, studies such as Gropp et al. (2020), Chen et al. (2020a), Lee et al. (2022), Horan et al. (2021), Nakagawa et al. (2023), Huh et al. (2023), Hahm et al. (2024) have provided a new perspective, showing that geometric constraints on representation spaces may also enable disentanglement. Typically, latent representations are encouraged to preserve key geometric features of the data distribution, such as (scaled) distances or angles between samples. Horan et al. (2021) even demonstrate that unsupervised disentanglement is always possible provided that the latent space is locally isometric to the data, further supporting the geometric desiderata. However, combining prior matching with these geometric aspects is challenging. In general, a mapping that perfectly aligns the data distribution with the prior while fully preserving the geometric features of interest may not exist. This leads to an inherent trade-off: Practitioners must carefully fine-tune regularization terms, either by altering prior matching to prioritize geometry preservation, or vice-versa.
In this work, we demonstrate how to effectively combine geometric desiderata with prior matching within the VAE framework, using optimal transport (OT) theory (Santambrogio, 2015, Peyré and Cuturi, 2019). By treating mappings from the data manifold to the latent space (encoders) or vice versa (decoders) as transport maps $T:\mathcal{X}\to\mathcal{Y}$, we can leverage the Gromov-{Monge, Wasserstein} paradigm (Sturm, 2020, Mémoli, 2011), which aligns two distributions by finding a mapping that minimizes the distortion between intra-domain cost functions defined on their supports. Specifically, we consider cost functions $c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})$ on $\mathcal{X}$ and $c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})$ on $\mathcal{Y}$ that encode geometric features such as scaled distances or angles. Consequently, the resulting mapping transforms one distribution onto the other while preserving these geometric features as much as possible.
Our Contribution: A novel OT-based approach to disentanglement through geometric considerations.
  • (i)
    We address the challenge of learning disentangled representations using geometric constraints by leveraging Gromov-Monge mappings between the data and prior distributions. Since fully preserving geometric features—such as (scaled) distances or angles between points—during the alignment of these two distributions is generally impossible, we aim to find an alignment that, instead, minimizes the distortion of these features, thereby preserving them as much as possible.
  • (ii)
    Inspired by (Uscidda and Cuturi, 2023), we introduce the Gromov-Monge Gap (GMG), a regularizer that measures how closely a map $T$ approximates a Gromov-Monge map for costs $c_{\mathcal{X}},c_{\mathcal{Y}}$. GMG measures whether $T$ transports distributions with minimal distortion w.r.t. $c_{\mathcal{X}},c_{\mathcal{Y}}$. We propose an efficient procedure to compute GMG and describe how to integrate it within the VAE framework.
  • (iii)
    We show that when $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ encode scaled distances or angles, the GMG and its finite-sample counterpart are weakly convex functions. In both cases, we precisely characterize the weak convexity constants and analyze their practical implications for practitioners.
  • (iv)
    Across four standard disentangled representation learning benchmarks, we show that incorporating geometry-preserving desiderata via the GMG significantly enhances disentanglement across various methods—from the standard $\beta$-VAE to the combination of $\beta$-TCVAE with HFS (Roth et al., 2023)—notably outperforming other methods that leverage geometric constraints.
@@ 2.1 Disentangled Representation Learning
Disentanglement has varying operational definitions. In this work, we follow the common understanding (Higgins et al., 2017, Locatello et al., 2019a; 2020, Träuble et al., 2021, Roth et al., 2023) where data $\mathbf{x}$ is generated by a process $p(\mathbf{x}|\mathbf{z})$ operating on ground-truth latent factors $\mathbf{z}\sim p(\mathbf{z})$, modeling underlying source of variations (s.a. object shape, color, background…). Given a dataset $\mathcal{D}=\{\mathbf{x}_{i}\}_{i=1}^{N}$, $\mathbf{x}_{i}\sim p_{\textrm{data}}$, unsupervised disentangled representation learning aims to find a mapping $e_{\phi}$ s.t. $e_{\phi}(\mathbf{x}_{i})\approx\mathbb{E}[\mathbf{z}|\mathbf{x}_{i}]$, up to element-wise transformations. Notably, this is to be achieved without prior information on $p(\mathbf{z})$ and $p(\mathbf{x}|\mathbf{z})$.
@@
Most unsupervised disentanglement methods operate on variational autoencoders (VAEs, Kingma and Welling (2014)), which define a generative model of the form $p_{\theta}(\mathbf{x},\mathbf{z})=p(\mathbf{z})p_{\theta}(\mathbf{x}|\mathbf{z})$. Here, $p_{\theta}(\mathbf{x}|\mathbf{z})$ is a product of exponential family distributions with parameters computed by a decoder $d_{\theta}(\mathbf{z})$. The latent prior $p(\mathbf{z})$ is usually chosen as a standard Gaussian $\mathcal{N}(\mathbf{z}|\mathbf{0}_{d},\mathbf{I}_{d})$, and the probabilistic encoder $q_{\phi}(\mathbf{z}|\mathbf{x})$ is implemented through neural networks $e_{\phi}(\mathbf{x}),\sigma_{\phi}(\mathbf{x})$ that predicts the latent parameters so that $q_{\phi}(\mathbf{z}|\mathbf{x})=\mathcal{N}(\mathbf{z}|e_{\phi}(\mathbf{x}),\sigma^{2}_{\phi}(\mathbf{x}))$. The $\beta$-VAE (Higgins et al., 2017) achieves disentanglement by minimizing
$\min_{\theta,\phi}\mathbb{E}_{\mathbf{x}\sim p_{\textrm{data}},\mathbf{z}\sim q_{\phi}(\mathbf{z}|\mathbf{x})}[\underbrace{-\log p_{\theta}(\mathbf{x}|\mathbf{z})}_{\textrm{(i) reconstruction}}+\underbrace{\beta D_{\text{KL}}(q_{\phi}(\mathbf{z}|\mathbf{x})||p(\mathbf{z}))}_{\textrm{(ii) prior matching}}],$ (1)
which enforces $\beta$-weighted prior matching on top of the reconstruction loss, assuming statistical factor independence (Roth et al., 2023). Several follow-ups refine latent prior matching through different losses or prior choices (Kim and Mnih, 2018, Chen et al., 2018, Kumar et al., 2018, Burgess et al., 2018, Rolinek et al., 2019, Moor et al., 2021, Balabin et al., 2024).
@@
Recent research suggests that disentanglement can arise by imposing geometric constraints on representation spaces (Horan et al., 2021). Most works (Karras et al., 2020, Huh et al., 2023) promote latent representations to preserve scaled distances between data points. Closest to our work, Lee et al. (2022) use a Jacobian-based regularization within a VAE, while Nakagawa et al. (2023) employ a distortion-based regularization, directly comparing pairwise squared Euclidean distances, within a Wasserstein AE (WAE, Tolstikhin et al. (2019)).
@@ 2.2 Quadratic Optimal Transport
OT (Peyré and Cuturi, 2019) involves transferring a probability distribution to another with inductive biases. When these distributions lie on incomparable domains, one relies on the Gromov-Monge (GM) and GW problems, also known as quadratic formulations of OT. Consider two compact $\mathcal{X}\subset\mathbb{R}^{d_{\mathcal{X}}}$, $\mathcal{Y}\subset\mathbb{R}^{d_{\mathcal{Y}}}$, each of them equipped with an intra-domain cost $c_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ and $c_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$. For technical reasons, we assume that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $-c_{\mathcal{X}}$ and $-c_{\mathcal{Y}}$) are CPD kernels (see Def. (A.1)). For $p\in\mathcal{P}(\mathcal{X})$ and $q\in\mathcal{P}(\mathcal{Y})$—two distributions supported on each domain—the GM problem (Mémoli and Needham, 2022) seeks a map $T:\mathcal{X}\to\mathcal{Y}$ that push-forwards $p$ onto $q$, while minimizing the distortion of the costs:
@@
When it exists, we call a solution $T^{\star}$ to (GMP) a Gromov-Monge map for costs $c_{\mathcal{X}},c_{\mathcal{Y}}$. However, this formulation is ill-suited for discrete distributions $p,q$, as the constraint set might be empty in that case. Replacing maps by coupling $\pi\in\Pi(p,q)$, i.e. distributions on $\mathcal{X}\times\mathcal{Y}$ with marginals $p$ and $q$, we obtain the GW problem (Mémoli, 2011, Sturm, 2020)
$\mathrm{GW}(p,q):=\min_{\pi\in\Pi(p,q)}\int_{(\mathcal{X}\times\mathcal{Y})^{2}}\tfrac{1}{2}|c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})|^{2}\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})\mathop{}\!\mathrm{d}\pi(\mathbf{x}^{\prime},\mathbf{y}^{\prime})\,.$ (GWP)
A solution $\pi^{\star}$ to (GWP) always exists, making $\mathrm{GW}(p,q)$ a well-defined quantity. It quantifies the minimal distortion of the geometries induced by $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ achievable when coupling $p$ and $q$.
@@
When both $p$ and $q$ are instantiated as samples, GW Prob. (GWP) translates to a quadratic assignment problem, whose objective can be regularized using entropy (Cuturi, 2013, Peyré et al., 2016). For empirical measures ${p}_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{\mathbf{x}_{i}}$, ${q}_{n}=\frac{1}{n}\sum_{j=1}^{n}\delta_{\mathbf{y}_{j}}$ and $\varepsilon\geq 0$, we set:
$\mathrm{GW}_{\varepsilon}({p}_{n},{q}_{n})\vcentcolon=\min_{\mathbf{P}\in U_{n}}\sum_{i,j,i^{\prime},j^{\prime}=1}^{n}(\mathbf{C}_{\mathcal{X}_{i,i^{\prime}}}-\mathbf{C}_{\mathcal{Y}_{j,j^{\prime}}})^{2}\,\mathbf{P}_{i,j}\mathbf{P}_{i^{\prime},j^{\prime}}-\varepsilon H(\mathbf{P})\,,$ (EGWP)
with $\mathbf{C}_{\mathcal{X}}=[c_{\mathcal{X}}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}})]_{i,i^{\prime}}$, $\mathbf{C}_{\mathcal{Y}}=[c_{\mathcal{Y}}(\mathbf{y}_{j},\mathbf{y}_{j^{\prime}})]_{j,j^{\prime}}\in\mathbb{R}^{n\times n}$, $U_{n}=\{\mathbf{P}\in\mathbb{R}^{n\times n}_{+},\mathbf{P}\mathbf{1}_{n}=\mathbf{P}^{T}\mathbf{1}_{n}=\tfrac{1}{n}\mathbf{1}_{n}\}$ and $H(\mathbf{P})=-\sum_{i,j=1}^{n}\mathbf{P}_{i,j}\log(\mathbf{P}_{i,j})$. As $\varepsilon\to 0$, we recover $\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}_{0}=\mathrm{GW}^{c_{\mathcal{X}},c_{\mathcal{Y}}}$. In addition to yielding greater convexity (Rioux et al., 2023) and better statistical (Zhang et al., 2023) and properties, entropic regularization improves computational performance. We can solve (EGWP) using a scheme that iterates the Sinkhorn algorithm (Peyré et al., 2016). This solver has $\mathcal{O}(n^{2})$ memory complexity. Its time complexity depends on $c_{\mathcal{X}},c_{\mathcal{Y}}$. For general $c_{\mathcal{X}},c_{\mathcal{Y}}$, it runs in $\mathcal{O}(n^{3})$. However, for the most common practical choices of $c_{\mathcal{X}}=c_{\mathcal{Y}}=\langle\cdot,\cdot\rangle$ or $c_{\mathcal{X}}=c_{\mathcal{Y}}=\|\cdot-\cdot\|_{2}^{2}$, it can also be reduced to $\mathcal{O}(n^{2}(d_{\mathcal{X}}+d_{\mathcal{Y}}))$, as detailed by Scetbon et al. (2022, §3 & Alg. 2).
Neural GW.
While for classical OT, numerous neural methods have been proposed (Makkuva et al., 2020, Korotin et al., 2022, Tong et al., 2023, Eyring et al., 2024), the GW setting has received less attention. Carrasco et al. (2024) propose an approach for solving GM with inner product costs, that involves a min-max-min optimization procedure. Klein et al. (2024) propose a framework to computeentropic GW couplings. Yet, these formulations are end-to-end, and not based on a regularizer that can be flexibly added to a loss function. Therefore, they cannot be combined with other models, as we do in this work with the $\beta$-VAE. In a concurrent work, Sotiropoulou and Alvarez-Melis (2024) introduced the Gromov-Monge gap. However, (i) they do not apply it for disentangled representational learning, and (ii) they do not investigate its properties, as we do in this work.